# Towards Scalability and Extensibility of Query Reformulation Modeling in E-commerce Search

### v1

@@ Abstract.
Customer behavioral data significantly impacts e-commerce search systems. However, in the case of less common queries, the associated behavioral data tends to be sparse and noisy, offering inadequate support to the search mechanism. To address this challenge, the concept of query reformulation has been introduced. It suggests that less common queries could utilize the behavior patterns of their popular counterparts with similar meanings. In Amazon product search, query reformulation has displayed its effectiveness in improving search relevance and bolstering overall revenue. Nonetheless, adapting this method for smaller or emerging businesses operating in regions with lower traffic and complex multilingual settings poses the challenge in terms of scalability and extensibility. This study focuses on overcoming this challenge by constructing a query reformulation solution capable of functioning effectively, even when faced with limited training data, in terms of quality and scale, along with relatively complex linguistic characteristics. In this paper we provide an overview of the solution implemented within Amazon product search infrastructure, which encompasses a range of elements, including refining the data mining process, redefining model training objectives, and reshaping training strategies. The effectiveness of the proposed solution is validated through online A/B testing on search ranking and Ads matching. Notably, employing the proposed solution in search ranking resulted in 0.14% and 0.29% increase in overall revenue in Japanese and Hindi cases, respectively, and a 0.08% incremental gain in the English case compared to the legacy implementation; while in search Ads matching led to a 0.36% increase in Ads revenue in the Japanese case.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: 30th SIGKDD Conference on Knowledge Discovery and Data Mining - Applied Data Science Track; August 25–29, 2024; Barcelona, Spainisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Information retrievalccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
In the realm of E-commerce, we consider the robustness of a QR solution from two key angles: scalability and extensibility. Scalability pertains to the size and quality of data required to support the effective model training. It is widely recognized that having a sufficient amount of high-quality training data can significantly ease the intricacies of the training process and reduce system complexity. However, for smaller markets or niche business fields, customer behavior signals often tend to be both sparse and noisy, making it challenging to construct a high-quality training dataset. Extensibility decides whether the QR approach can be applied across a broad spectrum of linguistic scenarios, including those with more intricate text patterns (take Japanese, for instance; unlike Indo-European languages, it comprises a mix of ideographic and phonetic characters, resulting in various script types, such as Kanji, Hiragana, Katakana, and Romaji, that may convey the same semantic meaning but are used in vastly different ways in diverse contexts). As such, our central focus in this paper is understanding how to build a QR solution capable of functioning effectively, even when faced with limited training data, in terms of quality and scale, along with relatively complex linguistic characteristics.
@@
The primary challenge in achieving robustness of QR stems from its reliance on customer behavior data - the development of the QR framework is susceptible to behavioral noise caused by uneven query exposure to customers (Szpektor et al., 2011), evident in the form of false positive/negative query-query pairs within the training/evaluation data extracted from historical customer behavior. This noise are particularly prevalent with cold-start queries and those with limited association to related products due to system defects, which are exactly part of QR’s target. In light of this challenge, we have refined core concepts and components of the legacy reformulation modeling (Huang et al., 2023). Our objective is to extend the benefit of the QR framework to smaller, non-English markets in addition to expand the benefit in major English markets. This paper presents a comprehensive overview of these improvements. Specifically, we encapsulate the finding and contribution of this study as follows,
  • We introduce novel measurements for query-query relationship mining with behavioral data, specifically designed to reduce biases stemming from false positives during downstream model training.
  • We refine model training objectives by introducing a sample importance factor, and also integrate hard negative mining into the model training pipeline to improve recall and precision, respectively at the retrieval and re-ranking stage of QR.
@@
  • Through comprehensive offline analyses, we illustrate the proposed approaches effectively mitigate behavioral noise, thereby enhancing offline recall@100 by 35.2% and NDCG@3 in the hard negative scenario by 94.4% compared with the legacy QR modeling. Furthermore, results from human audits indicate that these approaches markedly strengthen the model’s ability to discern between relevant and challenging irrelevant query pairs, resulting in a noteworthy improvement in classification AUC from 0.51 to 0.79.
@@ 2. Related work
Query reformulation pipeline With the advent of transformer-based text encoders, there has been a paradigm shift toward leveraging the semantic information encoded within queries for rewrite tasks (Chen et al., 2020; Cho et al., 2021), employing concepts like semantic textual similarity and dense text retrieval. This entails generating representative query embeddings, enabling the establishment of query-to-query relationships on a large scale in a computationally efficient manner. While there has been a notable effort to develop QR modeling within E-commerce applications (Li et al., 2022; Qiu et al., 2021), these endeavors did not specifically focus on enhancing the behavioral signal of less common queries, and thus, the design of these models was not tailored accordingly. A previous work (Huang et al., 2023) bridges this gap for Amazon product search with a query reformulation pipeline, comprising two components: 1) A Query-to-Query retrieval model, designed to identify a multitude of candidate queries that share a similar shopping intent with a given query; 2) A Query-to-Query re-ranking model, which refines the selection process by choosing the top queries with the highest semantic similarity from the pool of candidates. In this work, we focus on exploring the robustness issue in the application of QR methodologies, and more specifically, refining the design and implementation of this pipeline to benefit smaller or emerging businesses in regions characterized by multilingual or complex language contexts.
@@ 3. Problem formulation
In this section, we present a comprehensive formulation of the query reformulation problem within E-commerce search tasks. We begin with a set of search queries, denoted as $Q$, collected over a specific time frame. We divide this set into distinct subsets: $Q=Q_{r}+Q_{i}$, where $Q_{r}$ denotes behavior-rich queries, $Q_{i}$ denotes behavior-impoverished queries.
For the Query Reformulation (QR) problem, our objective is to establish mappings for each behavior-impoverished query, $q_{i}\in Q_{i}$, such that $q_{i}:\mapsto\{q_{i1},...,q_{iN}\}$, where $q_{ik}\in Q_{r}$, and each query mapping $(q_{i},q_{ik})$ is considered to be semantically relevant. Given the absence of a universal gold standard for defining the semantic relevance of query pairs, we also consider a sub-problem to derive a real-world relevance estimation on the query pair space $Q_{i}\times Q_{r}\mapsto s$, where $s$ denotes a relevance score.
@@ 4. Method
In this section, we elaborate on the enhancements introduced to the training of models mentioned above. Our refinements encompass the data mining process to generate training dataset, the model training objective, and an integration of Approximate Nearest Neighbor Negative Contrastive Estimation (ANCE) (Sohn, 2016) into both the retrieval and re-ranking model training processes.
@@
4.1. Mining query-query training data
The training of bi-encoder and cross-encoder models relies on positive and negative query-query pairs derived from mining customer behavior data (i.e., search log data). This subsection introduce a data mining strategy designed to mitigate false positive cases. The initial step involves extracting the behavior data (i.e., products and their corresponding purchase counts) of each query $q_{i}$, denoted as:
@@ 4.1. Mining query-query training data
where $p_{ik}$ represent a product which has been purchased under query $q_{i}$ and n signifies the corresponding number of purchase. This dataset exclusively consists of queries from $Q_{r}$ as other queries lack a sufficiently strong signal to form positive and negative labels. The behavior data is further refined as $B_{filtered}(q_{i})=\{p_{ik}|n_{ik}\geq P\}$ to filter out the query-product pairs with less than P purchase to reduce noise in the dataset. Finally, the pair $(q_{i},q_{j})$ is retained if $|B_{filtered}(q_{i})\cap B_{filtered}(q_{j})|\geq C$ where ||$\cdot$|| ⋅ | represents the cardinality of a set.
The purpose of the QR mechanism is to facilitate the sharing of behavioral signals between queries. However, it is crucial to mitigate the inadvertent introduction of irrelevant query-product associations into downstream search applications. The key principle is to ensure that, for any query $q_{i}$, only $q_{ik}$ relevant to the shopping intent of $q_{i}$ is retained.
One approach to enforce this constraint in the models’ practice is to prevent the occurrence of ”false positives” in the model training data. For instance, the queries “furniture” and “dining table” share numerous associated products, and are usually considered related in many behavioral mining algorithms. However, borrowing behavior of the query “furniture” to the query “dining table” would inevitably relate products like “TV cabinet” to “dining table”. A prior study (Huang et al., 2023) introduced an advanced positive labeling approach that quantifies the similarity between $q_{1}$ and $q_{2}$ as
@@
with $PP(q)$ representing the set of products purchased under query $q$. Only the top X(i.e., a pre-defined hyper-parameter) pairs associated to each query according to the similarity are defined as positive.
In this paper, we refine the similarity measurement as follows.
  1. (1)
    Normalize the number of purchase in $B_{filtered}(q_{i})$ as $\hat{n_{ik}}=n_{ik}/\sum_{m=1}^{x}n_{im}$.
  2. (2)
    Calculate Jensen–Shannon divergence (JSD) between normalized purchased product count distribution, denoted as $JSD(dist(q_{1})||dist(q_{2})$, where $dist(q_{1})=\{\hat{n_{11}},\hat{n_{12}},...,\hat{n_{1x}}\}$, $dist(q_{2})=\{\hat{n_{21}},\hat{n_{22}},...,\hat{n_{2y}}\}$.
The improved measurement alleviates a common issue where a pair might receive a high similarity score despite having significantly different distributions in their product purchase counts (as shown in Figure 1(a)). Under this refined measurement, only pairs that share not only the same set of purchased products but also a similar distribution can achieve a high similarity score (as shown in Figure 1(b)). To illustrate the impact of this enhancement, we provide examples of the top-1 similar query pairs $(q_{i},q_{i1})$ under both the previous and the new measurement methods in Table 1. It becomes evident that the new measurement yields more suitable query pairs and eliminates undesirable cases, such as ”google pixel watch” and ”pixel watch case.”
@@
$q_{i}$
Translation
$q_{i1}$ (previous)
Translation
$q_{i1}$ (proposed)
Translation
google pixel watch
N/A
pixel watch ケース
pixel watch case
pixel watch
N/A
usb bluetooth 5.0
N/A
bluetooth tplink
N/A
tp-link bluetooth usb bluetooth 5.0
N/A
pdp コントローラー
pdp controller
pdp
N/A
pdp コントローラー switch
pdp controller switch
n極対応 延長コード
n-pole compatible extension cord
n極対応
Compatible with n-pole
延長コード n極対応
n-pole compatible extension cord
To further mitigate the occurrence of false positives, we take into account the asymmetrical nature of $(q_{i},q_{ik})$ pairs in the training data. It is noteworthy that $(q_{i},q_{ik})$ forms a suitable pair in the context of QR does not necessarily imply that $(q_{ik},q_{i})$ is an appropriate pair. In the KNN search-based retrieval task, it is infeasible to finetune the model to differentiate between the positions of queries within a pair, as the embedding calculation only takes one query at a time as input. However, in the re-ranking task, the cross-encoder model can be trained in a position-aware mode. Therefore, we use separate similarity measures for retrieval and re-ranking tasks to fully exploit the inductive bias of each model and ultimately address relevance defects. First, JSD is used for retrieval as a symmetric optimization objective would stabilize the bi-encoder training process. Second, consider the calculation
@@
4.2. Query normalization and training data augmentation
In the realm of E-commerce, the volume of behavioral data for queries and products within the same category can fluctuate significantly, with differences of up to tenfold across various countries and languages. Successful QR practice in Amazon product search have traditionally leveraged an extensive dataset, encompassing over one billion query-query pairs derived from approximately thirty million unique queries. However, when dealing with countries with a small traffic and featuring a unique language, the availability of data can be notably limited. To address this challenge, we advocate the use of query normalization-based data augmentation in mining query-query data, considering the scale of available data required for effective model training.
@@ 4.2. Query normalization and training data augmentation
In developing our QR framework for non-English and low-traffic markets, we employ a text-normalization pipeline of preprocessing steps, including stop word removal, stemming, script type normalization (e.g., translation between different spellings of the same word, as illustrated by ”こども” to ”子供” in the Japanese case), tokenization (e.g., converting ”子供マスク不織布” to ”子供_マスク_不織布”), and token sorting (e.g., reordering ”子供_マスク_不織布” to ”マスク_子供_不織布”). It is imperative to note that this process excludes the handling of identified named entities, such as brands or media titles, in order to prevent any inadvertent introduction of defects.
Recall that in the construction of the query-query dataset, our approach involves two filtering steps aimed at reducing noise. These steps exclude query-product pairs with a low number of purchases and query-query pairs with a limited number of co-purchased products. As a consequence, a considerable number of relevant query-query pairs are eliminated due to the sparsity of behavioral signals. The introduction of query normalization offers a significant mitigation to this issue. In this study, it substantially increases the count of query-query pairs from about 9 million to about 348 million. Furthermore, it also expands the number of unique queries with low request traffic (i.e., contributing to the last one-third of traffic) covered by the behavioral dataset, elevating it from about 166 thousand to more than 3 million. Please note that this number is derived from the entire query set rather than $Q_{r}$.
@@
Furthermore, it has come to our attention that the similarity scores derived from behavioral data between each member within a group of queries sharing the same shopping intent but slightly different representations can exhibit significant disparities. We illustrate this phenomenon in Table 2, where we present the similarity scores computed between the aggregated behavior of a group of queries all centered around the same shopping intent (in this case, masks for children made of non-woven fabric) and the individual behavior of each query within the group. The queries in this group only differ in terms of tokenization (i.e., the presence or absence of space) and token order, but the similarity scores range from 0.18 to 0.71 and 0.22 to 0.62, respectively. Hence, we contend that clustering query behavior through normalization is a necessary step to acquire similarity scores that more accurately reflect the semantic distinctions between queries.
Table 2. An example in Japanese of query similarity disparity within group. In both groups, the queries convey the same meaning, referring to ”masks for children made of non-woven fabric”, but are in different representations with respect to script-type and tokenization (e.g., the text in gray boxes all refers to ”for children”). The normalized representation of the queries within each group is identical.
Group 1
Query Score
子供_マスク_不織布 0.18
マスク_不織布_子供 0.71
マスク_子ども_不織布 0.44
マスク_不織布_こども 0.30
Group 2
Query Score
子供用マスク_不織布 0.62
マスクこども用不織布 0.29
ことも用マスク不織布 0.22
マスク_子供用_不織布 0.40
4.3. Model training with ANCE mining

### v2

@@ Abstract.
Customer behavioral data significantly impacts e-commerce search systems. However, in the case of less common queries, the associated behavioral data tends to be sparse and noisy, offering inadequate support to the search mechanism. To address this challenge, the concept of query reformulation has been introduced. It suggests that less common queries could utilize the behavior patterns of their popular counterparts with similar meanings. In Amazon product search, query reformulation has displayed its effectiveness in improving search relevance and bolstering overall revenue. Nonetheless, adapting this method for smaller or emerging businesses operating in regions with lower traffic and complex multilingual settings poses the challenge in terms of scalability and extensibility. This study focuses on overcoming this challenge by constructing a query reformulation solution capable of functioning effectively, even when faced with limited training data, in terms of quality and scale, along with relatively complex linguistic characteristics. In this paper we provide an overview of the solution implemented within Amazon product search infrastructure, which encompasses a range of elements, including refining the data mining process, redefining model training objectives, and reshaping training strategies. The effectiveness of the proposed solution is validated through online A/B testing on search ranking and Ads matching. Notably, employing the proposed solution in search ranking resulted in 0.14% and 0.29% increase in overall revenue in Japanese and Hindi cases, respectively, and a 0.08% incremental gain in the English case compared to the legacy implementation; while in search Ads matching led to a 0.36% increase in Ads revenue in the Japanese case. The proposed solution has been deployed in Amazon product search.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, Idaho, USAisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Information retrievalccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
In the realm of E-commerce, we consider the robustness of a QR solution from two key angles: scalability and extensibility. Scalability pertains to the size and quality of data required to support the effective model training. It is widely recognized that having a sufficient amount of high-quality training data can significantly ease the intricacies of the training process and reduce system complexity. However, for smaller markets or niche business fields, historical customer engagement often tend to be both sparse and noisy, making it challenging to construct a high-quality training dataset. Extensibility decides whether the QR approach can be applied across a broad spectrum of linguistic scenarios, including those with more intricate text patterns.111take Japanese, for instance; unlike Indo-European languages, it comprises a mix of ideographic and phonetic characters, resulting in various script types, such as Kanji, Hiragana, Katakana, and Romaji, that may convey the same semantic meaning but are used in vastly different ways in diverse contexts As such, our central focus in this paper is understanding how to build a QR solution capable of functioning effectively, even when faced with limited training data, in terms of quality and scale, along with relatively complex linguistic characteristics.
@@
The primary challenge in achieving robustness of QR stems from its reliance on historical customer engagement - the development of the QR framework is susceptible to noise in training/evaluation data caused by uneven query exposure to customers (Szpektor et al., 2011), evident in the form of false positive/negative query-query pairs. This noise are particularly prevalent with cold-start queries and those with limited association to related products due to system defects, which are exactly part of QR’s target. In light of this challenge, we have refined core concepts and components of the legacy reformulation modeling (Huang et al., 2023). Our objective is to extend the benefit of the QR framework to smaller, non-English markets in addition to expand the benefit in major English markets. This paper presents a comprehensive overview of these improvements. Specifically, we encapsulate the finding and contribution of this study as follows,
  • We refine model training objectives by introducing a sample importance mining approach, which enables better data utilization in data resource-limited scenario. We also integrate hard negative mining into the re-ranking model training through a learn-from-teacher strategy, which is demonstrated critical to ensure the model’s performance.
@@
  • Through comprehensive offline analyses, we illustrate the proposed approaches enhance offline recall@100 by 35.2% and NDCG@3 in the hard negative scenario by 94.4% compared with the legacy QR modeling. Furthermore, results from human audits indicate that these approaches markedly strengthen the model’s ability to discern between relevant and challenging irrelevant query pairs, resulting in a noteworthy improvement in classification AUC from 0.51 to 0.79.
@@ 2. Related work
Query reformulation pipeline With the advent of transformer-based text encoders, there has been a paradigm shift toward leveraging the semantic information encoded within queries for rewrite tasks (Chen et al., 2020; Cho et al., 2021), employing concepts like semantic textual similarity and dense text retrieval. This entails generating representative query embeddings, enabling the establishment of query-to-query relationships on a large scale in a computationally efficient manner. While there has been a notable effort to develop QR modeling within E-commerce applications (Li et al., 2022; Qiu et al., 2021), these endeavors did not specifically focus on enhancing the behavioral signal of less common queries, and thus, the design of these models was not tailored accordingly. A previous work (Huang et al., 2023) bridges this gap for Amazon product search with a query reformulation pipeline. In this work, we focus on exploring the robustness issue in the application of QR methodologies, and more specifically, refining the design and implementation of this pipeline to benefit smaller or emerging businesses in regions characterized by multilingual or complex language contexts.
@@ 3. Problem formulation
In this section, we present a formulation of the query reformulation problem within the behavior enrichment task. We begin with a set of search queries, denoted as $Q$, collected over a specific time frame. We divide this set into distinct subsets: $Q=Q_{r}+Q_{i}$, where $Q_{r}$ denotes behavior-rich queries, $Q_{i}$ denotes behavior-impoverished queries.
For the Query Reformulation (QR) problem, our objective is to establish mappings for each behavior-impoverished query, $q_{i}\in Q_{i}$, such that $q_{i}:\mapsto\{q_{i1},...,q_{iN}\}$, where $q_{ik}\in Q_{r}$, and each query mapping $(q_{i},q_{ik})$ is considered to be semantically relevant.
@@ 4. Method
In this section, we elaborate on the enhancements introduced to the training of models mentioned above. Our refinements encompass the model training objective, the data mining process to generate signals for training, and an integration of Approximate Nearest Neighbor Negative Contrastive Estimation (ANCE) (Sohn, 2016) into the re-ranking model training processes.
@@
4.1. Retrieval model training schema
We use $s_{BE}(q_{i},q_{j})$ to represent the cosine similarity between L2 normalized embedding of $q_{i}$ and $q_{j}$. For retrieval task, we define the training objective under the contrastive learning with in-batch negative framework as
(1) $\mathcal{L}_{\mathrm{retrieval}}=-1/N\cdot\Sigma_{i,k}I(q_{i},q_{ik})\cdot\frac{\exp{s_{BE}(q_{i},q_{ik})/\tau}}{\Sigma_{j\in\mathrm{batch}}\exp{s_{BE}(q_{i},q_{j})/\tau}}$
where $N$ is the number of positive query pairs, $\tau$ is a temperature parameter. Specifically, we add a sample importance term $I(q_{i},q_{ik})$ to the infoNCE (Oord et al., 2018) (NT-XENT (Sohn, 2016)) objective. We first explain the necessity of adding this term and leave the detailed calculation to the next subection.
RQ1: Why is sample importance critical for the QR solution in the resource-limited scenario?
Fact1.1: In the context of QR, the positives are derived from behavioral data, which are prone to being false positives. For instance, the queries “furniture” and “dining table” share numerous associated products, and are usually considered related in many behavioral mining algorithms. However, borrowing behavior of the query “furniture” to the query “dining table” would inevitably relate products like “TV cabinet” to “dining table”, and consequently, significantly compromises search relevance.
Fact1.2: In resource-limited scenario, simply remove semi-relevant query pairs could harm the training efficacy. As we will show later in the experiment section, in a setting where we only kept top 30% percent of positive query pairs to ensure a high-quality training dataset, the experiment succeeded in the English case where the training data is abundant, but failed for languages with a relatively insufficiently large dataset. By introducing sample importance, on one hand, in the presence of abundant training data, the impact of less pertinent pairs is mitigated, allowing for more focused model learning. On the other hand, in scenarios with limited supervised resources, pairs that may not be strictly relevant-yet still bear some relevance-can still make a valuable contribution to the model’s training.
4.2. Sample importance through behavioral mining
This subsection introduce a data mining strategy to derive sample importance for each positive query-query pair. The initial step involves extracting the behavior data (i.e., products and their corresponding purchase counts) of each query $q_{i}$, denoted as:
@@ 4.2. Sample importance through behavioral mining
where $p_{ik}$ represent a product which has been purchased under query $q_{i}$ and $n$ signifies the corresponding number of purchase.
We define the query-query relevance measurement as follows.
  1. (1)
    Normalize the number of purchase in $B(q_{i})$ as $\hat{n_{ik}}=n_{ik}/\sum_{m=1}^{x}n_{im}$.
  2. (2)
    Calculate Jensen–Shannon divergence (JSD) between normalized purchased product count distributions $\{\hat{n_{11}},\hat{n_{12}},...,\hat{n_{1x}}\}$ and $\{\hat{n_{21}},\hat{n_{22}},...,\hat{n_{2y}}\}$.
This definition is an improvement based on a prior study(Huang et al., 2023), where the query relevance is defined as
@@
with $PP(q)$ representing the set of products purchased under query $q$.
The proposed measurement alleviates the issue that a pair might receive a high relevance score despite having significantly different distributions in their product purchase counts (as shown in Figure 2(a)). Under the proposed measurement, only pairs that share not only the same set of purchased products but also a similar distribution can achieve a high similarity score (as shown in Figure 2(b)).
To illustrate the impact of this enhancement, we provide examples of the top-1 similar query pairs $(q_{i},q_{i1})$ under both measurements method in Table 1. It becomes evident that the proposed measurement yields more suitable query pairs and eliminates undesirable cases, such as ”google pixel watch” and ”pixel watch case.”
@@
$q_{i}$ $q_{i1}$ (previous) $q_{i1}$ (proposed)
google pixel watch pixel watch case pixel watch
usb bluetooth 5.0 bluetooth tplink tp-link bluetooth usb bluetooth 5.0
pdp controller pdp pdp controller switch
4.3. Re-ranking model training schema
RQ2: What benefit does the re-ranking stage bring under the QR context?
Fact2: We take into account the asymmetrical nature of $(q_{i},q_{ik})$ pairs in the training data. It is noteworthy that $(q_{i},q_{ik})$ forms a suitable pair in the context of QR does not necessarily imply that $(q_{ik},q_{i})$ is also an appropriate pair. In the KNN search-based retrieval task, it is infeasible to finetune the model to differentiate between the positions of queries within a pair, as the embedding calculation only takes one query at a time as input. However, in the re-ranking task, the cross-encoder model can be trained in a position-aware mode, which enables the re-ranking stage to further mitigate false-positives in the end-to-end data generation process.
Specifically, we use separate similarity measures for retrieval and re-ranking tasks to fully exploit the inductive bias of each model and ultimately address relevance defects. Consider the calculation
@@ 4.3. Re-ranking model training schema
4.4. Query normalization for denoising
{CJK}
@@ 4.4. Query normalization for denoising
UTF8min In developing our QR framework for non-English and low-traffic markets, we employ a text-normalization pipeline of preprocessing steps, including stop word removal, stemming, script type normalization (e.g., translation between different spellings of the same word, as illustrated by ”こども” to ”子供” in the Japanese case), tokenization (e.g., converting ”子供マスク不織布” to ”子供_マスク_不織布”), and token sorting (e.g., reordering ”子供_マスク_不織布” to ”マスク_子供_不織布”). It is imperative to note that this process excludes the handling of identified named entities, such as brands or media titles, in order to prevent any inadvertent introduction of defects.
RQ3: Why is query normalization necessary?
@@
Fact3.1 The data quality perspective: It has come to our attention that the similarity scores (i.e., the sample importance used for model training) derived from behavioral data is noisy. Within a group of queries with the same shopping intent but slightly different representations, the similarity score between each pair of queries could have a discrepancy. We illustrate this phenomenon in Table 2, where we present a group of queries all centered around the same shopping intent. We show the similarity scores computed between the aggregated behavior of group and the individual behavior of each query within the group. The queries in this group only differ in terms of tokenization (i.e., the presence or absence of space) and token order, but the similarity scores range from 0.18 to 0.71. Hence, we contend that clustering query behavior through normalization is a necessary step to acquire similarity scores that more accurately reflect the semantic distinctions between queries.
Fact3.2 The data quantity perspective: In the construction of positive query-query pair training dataset, we use a filtering step aimed at reducing noise. This step exclude query-product pairs with a low number of purchases. As a consequence, a considerable number of relevant query-query pairs are eliminated due to the sparsity of behavioral signals. The introduction of query normalization offers a significant mitigation to this issue. Specifically, by grouping queries that are identical after normalization and combining their behavioral signals, we can enable a greater number of query-product pairs to surpass the filtering threshold. In this study, it substantially increases the count of query-query pairs from about 9 million to about 348 million. Furthermore, it also expands the number of unique queries with low request traffic (i.e., contributing to the last one-third of traffic) covered by the behavioral dataset, elevating it from about 166 thousand to more than 3 million. (Please note that this number is derived from the entire query set rather than $Q_{r}$.)
{CJK}
UTF8min Query Translation Relevance Score 子供_マスク_不織布 non-woven 0.18 マスク_不織布_子供 fabric 0.71 マスク_子ども_不織布 masks for 0.44 マスク_不織布_こども children 0.30
Table 2. An example in Japanese of query similarity disparity within a group. The queries have the same normalized representation but have varied relevance score.
4.5. Model training with hard negative mining